# Elucidated Rolling Diffusion Models for Probabilistic Weather Forecasting

### v1

@@ 4.1 Training Objective
The preconditioning functions $c_{\text{skip}},c_{\text{out}},c_{\text{in}},c_{\text{noise}}$ are adapted from EDM and applied per-snapshot based on the corresponding $\bar{\sigma}_{w}(t)$ in $\bm{\bar{\sigma}}({t})$. The denoiser network is trained to predict the clean window ${\bm{y}}_{1:W}$ from its noisy version, $\bar{\bm{x}}\sim\mathcal{N}({\bm{y}}_{1:W},\bm{\bar{\sigma}}(t)^{2}\mathbf{I})$. During training, we randomize $t~{}\sim U([0,1))$ such that the network learns to deal with the full extent of the segments of Fig. 2.
@@
where we shorten $\bm{\sigma}:=\bm{\bar{\sigma}}(t)$. The full training algorithm is described in Algorithm 1.
@@
5:     $\bar{\bm{x}}\leftarrow{\bm{y}}+\bm{\sigma}\cdot\bm{\epsilon}$triangleright Add rolling noise to data
@@
In both training and sampling algorithms, we draw i.i.d. Gaussian noise $\bm{\epsilon}_{w}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{D})$ for each snapshot $w$ independently. Recent studies show that temporally correlated noise improves video diffusion [8, 15, 1, 30]. Such priors are orthogonal to our sliding-window framework and can replace the i.i.d. draws in Algorithms 12. In our experiments, we use the progressive noise model with $\alpha=1$ proposed by Ge et al. [15], which we found to slightly improve long-range forecasts (see Appendix E). More details are provided in Section C.4.
@@ 4.2 Sampling
The trained denoiser, $D_{\theta}$, can be used to estimate the score function, $\log p(\bar{\bm{x}};\bm{\bar{\sigma}}({t}))$, in Eq. 2 with $(D_{\theta}(\bar{\bm{x}},\bm{\bar{\sigma}}({t}))-\bar{\bm{x}})/\bm{\bar{\sigma}}^{2}({t})$ for any ${t}\in[0,1)$ [23]. Based on this identity and the described probability flow ODE intuition, a mathematical formulation of our proposed sampling algorithm is provided in Algorithm 2. For simplicity, we omit the second-order Heun step and sampling stochasticity, which are detailed in Section C.2. Note, however, that in our main experiments, we always use the deterministic Heun sampler.
@@
The backward ODE requires a noisy window $\bar{\bm{x}}_{1:W}$ whose clean latent ${\bm{y}}_{1:W}$ is unknown. Ruhe et al. [46] addresses this issue by co-training the rolling model to generate $\bar{\bm{x}}_{1:W}$ from clean context snapshots and a sequence of pure noise. This approach introduces extra hyperparameters and diverts the denoiser’s capacity between two distinct denoising tasks. As an alternative, we propose drawing $\hat{\bm{y}}_{1:W}$ from an external forecaster $p_{\theta}({\bm{y}}_{1:W}|{\bm{y}}_{0})$. We then sample $\bar{\bm{x}}_{w}\sim\mathcal{N}\!\bigl{(}\hat{\bm{y}}_{w},\,\bar{\sigma}^{2}_{w}(0)\mathbf{I}_{D}\bigr{)}$ for $w=1,\dots,W$, from where we can start our sliding-window sampling algorithm. This choice reuses mature short-range models and injects schedule-matched uncertainty; any existing forecaster (diffusion-based or not) can supply $\hat{\bm{y}}_{1:W}$. In our experiments, we rely on the EDM baselines for initializing ERDM. We ablate this choice in Section E.1.
@@ 4.3 Denoiser architecture for temporal dynamics
Our proposed diffusion model operates on temporal sequences, necessitating an architecture capable of capturing temporal dependencies. Naive adaptations of 2D architectures, such as stacking the time dimension into channels, disrupt the inherent temporal structure and led to suboptimal results in our experiments (see Section 5.4). Alternatively, extending to 3D convolutions, while preserving temporality, often incurs significant computational overhead and potentially higher sample complexity. Instead, we adopt a prevalent strategy from latent video diffusion models [5]: augmenting a well-established 2D U-Net architecture with explicit temporal processing layers. Specifically, we integrate causal temporal attention layers into the 2D ADM U-Net [12] used in EDM, positioning before each down- and up-sampling block. These temporal layers also incorporate noise level information via a mechanism analogous to the adaptive layer normalization used in the 2D U-Net blocks. Figure 3 illustrates this interleaved modular design. This modular design could facilitate static pre-training of the 2D backbone, yet we observed that end-to-end training on the sequential data was more effective for our tasks (see Appendix E). While we found the proposed architecture to be key for generating good forecasts, it is slower and has higher memory needs than a 2D variant.
@@
2:$\Delta t\leftarrow 1/N$triangleright Infer step size from desired number of steps per snapshot
3:$t_{\rm{cur}}\leftarrow 0;\;\mathcal{S}\leftarrow\emptyset$triangleright Initialize global diffusion time and empty generated sequence
4:sample $\bar{\bm{x}}_{\rm{cur}}\sim\mathcal{N}(\hat{{\bm{y}}}_{1:W},\bm{\bar{\sigma}}(t_{\rm{cur}})^{2}\mathbf{I}_{W\times D})$triangleright Initialize window with snapshot-dependent rolling noise
@@
6:    $t_{\rm{next}}\leftarrow t_{\rm{cur}}+\Delta t$triangleright Global diffusion time after denoising
7:    $\bm{\sigma}_{\rm{cur}}\leftarrow\bm{\bar{\sigma}}(t_{\rm{cur}})$triangleright Current noise levels
8:    $\bm{\sigma}_{\rm{next}}\leftarrow\bm{\bar{\sigma}}(t_{\rm{next}})$triangleright Noise levels at the end of this iteration
9:    $\hat{{\bm{y}}}\leftarrow D_{\theta}(\bar{\bm{x}}_{\rm{cur}},\bm{\sigma}_{\rm{cur}})$triangleright Denoise sequence
10:    $\bm{d}\leftarrow(\bar{\bm{x}}_{\rm{cur}}-\hat{{\bm{y}}})/\bm{\sigma}_{\rm{cur}}$triangleright Evaluate $\mathrm{d}\bar{\bm{x}}/\mathrm{d}{t}$ at $t_{\rm{cur}}$
11:    $\bar{\bm{x}}_{\rm{next}}\leftarrow\bar{\bm{x}}_{\rm{cur}}+(\bm{\sigma}_{\rm{next}}-\bm{\sigma}_{\rm{cur}})\bm{d}$triangleright Euler step from $t_{\rm{cur}}$ to $t_{\rm{next}}$
12:    $n_{\rm{clean}}\leftarrow\lfloor t_{\rm{next}}\rfloor$triangleright Infer finished snapshots. If 00, the following lines are a no-op
@@
15:    $\bar{\bm{x}}_{\rm{cur}}\leftarrow[\bar{\bm{x}}_{\rm{next}},{\bm{x}}_{\mathrm{new}}]$triangleright Concatenate fresh noisy snapshots to futuremost
16:    $t_{\rm{cur}}\leftarrow t_{\rm{next}}-n_{\rm{clean}}$triangleright Re-adjust global diffusion ‘time’ to be in $[0,1)$
@@ 5.1 Evaluation and Metrics
Due to the importance of probabilistic forecasts and uncertainty quantification, we focus on two key ensemble-based probabilistic metrics: Continuous Ranked Probability Score (CRPS) [34], and the spread-skill ratio (SSR) based on $M$-member ensembles, where $M=50$ for Navier-Stokes and $M=10$ for ERA5. The CRPS is a proper scoring rule commonly used to evaluate probabilistic forecasts [16]. The spread-skill ratio is defined as the ratio of the square root of the ensemble variance to the corresponding ensemble-mean RMSE. It serves as a measure of the reliability of the ensemble, where values smaller than 1 indicate underdispersion, and larger values overdispersion [14, 43].
@@ 5.2 Navier-Stokes fluid dynamics
We use the Navier-Stokes fluid dynamics benchmark from [38], defined on a $221\times 42$ grid. Each simulation features four randomly placed circular obstacles influencing the flow, with fluid viscosity set to $1\times 10^{-3}$. The dataset comprises $x$ and $y$ velocities and pressure fields. All models receive boundary conditions and obstacle masks as auxiliary inputs. For testing, models predict a 64-timestep trajectory from a single initial snapshot.
@@
We benchmark ERDM against DYffusion [47], the current state-of-the-art method on this dataset. To ensure a fair comparison, we retrained DYffusion using our experimental setup, achieving an improvement of over 3×3$\times$3 × on its originally reported CRPS scores. Our primary focus, however, is to evaluate ERDM’s performance relative to common EDM-derived approaches for dynamics forecasting. Consequently, our key baseline is an EDM denoiser parameterized by $D_{\theta}({\bm{x}}_{1}\ldots,{\bm{x}}_{W};\sigma,{\bm{y}}_{0})$, where all ${\bm{x}}_{w}$ are corrupted according to the same $\sigma$. With $W=1$ we recover a next-step forecasting conditional EDM as in [42, 26]. We tried $W\in\{1,2,4,6\}$, and found $W=4$ to work best for EDM. All baselines are trained on three random seeds and share the same architecture as much as possible. For ERDM, we use $W=6,\sigma_{\text{min}}=0.002,\sigma_{\text{max}}=200,\rho=-10,P_{\text{mean}}=0.5,P_{\text{std}}=1.2,N=1.25$. Further details are provided in Section D.2.
@@ Results.
In Fig. 4, we benchmark ERDM against the baselines on forecasting the five Navier-Stokes test trajectories of length 64. While the EDM-based baselines start well in terms of CRPS, even beating ERDM for the very first few time steps, they exhibit a significantly higher error growth than ERDM. We attribute this to the explicit modeling of progressive uncertainty built into our model. At the end of the rollout, ERDM achieves a $50\%$ better CRPS than the best baseline, EDM with a window size of $W=4$. Similarly, DYffusion starts worse and catches up with the EDM $W=4$ baseline at the end of the rollout, but not with ERDM. In terms of calibration, we observe that ERDM consistently outperforms the EDM baselines, which are noticeably undercalibrated. We note that this is despite comprehensive tuning of these baselines, where we were only able to achieve calibration improvements at the significant expense of CRPS skill (not shown). MSE scores, qualitative results, and example generated videos are discussed in Section F.1.
@@ 5.3 ERA5 weather forecasting
We benchmark our model on medium-range weather forecasting on the ERA5 reanalysis dataset. We use the $1.5^{\circ}$ resolution of the data (a $240\times 121$ grid) provided by [43], with 69 prognostic (input and output) variables: Temperature (t), geopotential (z), specific humidity (q), and the u and v components of wind (u,v) over 13 pressure levels (numbers after the abbreviation refer to the level in hPa) as well as the four surface variables 2m temperature (2t), mean sea level pressure (mslp), and u and v components of winds at 10m (10u, 10v). We train our models on $12$-hourly data, as in [42], from 1979 to 2020 and evaluate on 64 2021 initial conditions at 00/12 UTC.
@@
Our primary baseline is a conditional next-step forecasting EDM, trained using a methodology consistent with ERDM’s development. To gauge the absolute performance of ERDM, we also benchmark it against two prominent external models: (1) IFS ENS: The European Centre for Medium-Range Weather Forecasts’ (ECMWF) operational, physics-based ensemble forecasting system [13]; (2) NeuralGCM ENS: A hybrid ML-physics stochastic model [25]. Official $1.5^{\circ}$-resolution forecasts for both external models were sourced from Weatherbench-2 [43] for all variables where available. The IFS ENS evaluation uses the same 2021 initial condition dates as our EDM and ERDM models, but is verified against operational analyses (as opposed to ERA5), as recommended by [43]. Due to the unavailability of NeuralGCM forecasts for 2021, we use its 2020 forecasts and evaluate them against the corresponding 2020 ERA5 data. Our conditional EDM baseline can be seen as a reproduction of GenCast [42], for which forecasts are only available at higher spatial resolutions and a limited set of variables, but using our experimental setup and neural architecture. Notably, training ERDM is considerably less computationally expensive than NeuralGCM ENS $1.5^{\circ}$ (GenCast $1^{\circ}$), which required 128 (32) v5 TPUs and 10 (3.5) days to train. In contrast, ERDM was trained on 4 H200 GPUs in 5 days only. Lastly, we also include scores for z500 and 2t from Graph-EFM [37], taken from their paper. These scores are provided for reference, noting that Graph-EFM is trained on 20 more years and evaluated in 2020. For ERDM, we use $W=6,\sigma_{\text{min}}=0.002,\sigma_{\text{max}}=500,\rho=-10,P_{\text{mean}}=2,P_{\text{std}}=1.2,N=2$. See Section D.2 for more experimental details.
Figure 6: Normalized spectral density of 14-day forecasts, averaged over high latitudes, $[60^{\circ},90^{\circ}]$. Spectra are divided by the target ERA5 reanalysis spectra. ERDM generates highly accurate spectra that match or slightly beat the physics-based model IFS ENS. NeuralGCM ENS underestimates energy at the mid to high frequencies. See Section F.3 for more variables and absolute spectra.

### v3

@@ 4.1 Training Objective
The preconditioning functions $c_{\text{skip}},c_{\text{out}},c_{\text{in}},c_{\text{noise}}$ are adapted from EDM and applied per-snapshot based on the corresponding $\bar{\sigma}_{w}(t)$ in $\bm{\bar{\sigma}}({t})$. The denoiser network is trained to predict the clean window ${\bm{y}}_{1:W}$ from its noisy version, $\bar{\bm{x}}\sim\mathcal{N}({\bm{y}}_{1:W},\bm{\bar{\sigma}}(t)^{2}\mathbf{I})$. During training, we randomize $t~\sim U([0,1))$ such that the network learns to deal with the full extent of the segments of Fig.˜2.
@@
where we shorten $\bm{\sigma}:=\bm{\bar{\sigma}}(t)$. The full training algorithm is described in Algorithm˜1.
@@
5:  $\bar{\bm{x}}\leftarrow{\bm{y}}+\bm{\sigma}\cdot\bm{\epsilon}$ $\triangleright$ Add rolling noise to data
@@
In both training and sampling algorithms, we draw i.i.d. Gaussian noise $\bm{\epsilon}_{w}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{D})$ for each snapshot $w$ independently. Recent studies show that temporally correlated noise improves video diffusion [chang2024warpednoise, ge2024noiseprior, andrae2025continuous, liu2025equivdm]. Such priors are orthogonal to our sliding-window framework and can replace the i.i.d. draws in Algorithms 12. In our experiments, we use the progressive noise model with $\alpha=1$ proposed by ge2024noiseprior, which we found to slightly improve long-range forecasts (see Appendix˜E). More details are provided in Section˜C.4.
@@ 4.2 Sampling
The trained denoiser, $D_{\theta}$, can be used to estimate the score function, $\log p(\bar{\bm{x}};\bm{\bar{\sigma}}({t}))$, in Eq.˜2 with $(D_{\theta}(\bar{\bm{x}},\bm{\bar{\sigma}}({t}))-\bar{\bm{x}})/\bm{\bar{\sigma}}^{2}({t})$ for any ${t}\in[0,1)$ [karras2022edm]. Based on this identity and the described probability flow ODE intuition, a mathematical formulation of our proposed sampling algorithm is provided in Algorithm˜2. For simplicity, we omit the second-order Heun step and sampling stochasticity, which are detailed in Section˜C.2. Note, however, that in our main experiments, we always use the deterministic Heun sampler.
@@
The backward ODE requires a noisy window $\bar{\bm{x}}_{1:W}$ whose clean latent ${\bm{y}}_{1:W}$ is unknown. ruhe2024rolling addresses this issue by co-training the rolling model to generate $\bar{\bm{x}}_{1:W}$ from clean context snapshots and a sequence of pure noise. This approach introduces extra hyperparameters and diverts the denoiser’s capacity between two distinct denoising tasks. As an alternative, we propose drawing $\hat{\bm{y}}_{1:W}$ from an external forecaster $p_{\theta}({\bm{y}}_{1:W}|{\bm{y}}_{0})$. We then sample $\bar{\bm{x}}_{w}\sim\mathcal{N}\!\bigl(\hat{\bm{y}}_{w},\,\bar{\sigma}^{2}_{w}(0)\mathbf{I}_{D}\bigr)$ for $w=1,\dots,W$, from where we can start our sliding-window sampling algorithm. This choice reuses mature short-range models and injects schedule-matched uncertainty; any existing forecaster (diffusion-based or not) can supply $\hat{\bm{y}}_{1:W}$. In our experiments, we rely on the EDM baselines for initializing ERDM. We ablate this choice in Section˜E.1.
@@ 4.3 Denoiser architecture for temporal dynamics
Our proposed diffusion model operates on temporal sequences, necessitating an architecture capable of capturing temporal dependencies. Naive adaptations of 2D architectures, such as stacking the time dimension into channels, disrupt the inherent temporal structure and led to suboptimal results in our experiments (see Section˜5.4). Alternatively, extending to 3D convolutions, while preserving temporality, often incurs significant computational overhead and potentially higher sample complexity. Instead, we adopt a prevalent strategy from latent video diffusion models [blattmann2023stablevideodiffusion]: augmenting a well-established 2D U-Net architecture with explicit temporal processing layers. Specifically, we integrate causal temporal attention layers into the 2D ADM U-Net [dhariwal2021diffusion] used in EDM, positioning before each down- and up-sampling block. These temporal layers also incorporate noise level information via a mechanism analogous to the adaptive layer normalization used in the 2D U-Net blocks. Figure 3 illustrates this interleaved modular design. This modular design could facilitate static pre-training of the 2D backbone, yet we observed that end-to-end training on the sequential data was more effective for our tasks (see Appendix˜E). While we found the proposed architecture to be key for generating good forecasts, it is slower and has higher memory needs than a 2D variant.
@@
[…]
4:sample $\bar{\bm{x}}_{\rm{cur}}\sim\mathcal{N}(\hat{{\bm{y}}}_{1:W},\bm{\bar{\sigma}}(t_{\rm{cur}})^{2}\mathbf{I}_{W\times D})$ $\triangleright$ Initialize window with snapshot-dependent rolling noise
@@
6:  $t_{\rm{next}}\leftarrow t_{\rm{cur}}+\Delta t$ $\triangleright$ Global diffusion time after denoising
7:  $\bm{\sigma}_{\rm{cur}}\leftarrow\bm{\bar{\sigma}}(t_{\rm{cur}})$ $\triangleright$ Current noise levels
8:  $\bm{\sigma}_{\rm{next}}\leftarrow\bm{\bar{\sigma}}(t_{\rm{next}})$ $\triangleright$ Noise levels at the end of this iteration
9:  $\hat{{\bm{y}}}\leftarrow D_{\theta}(\bar{\bm{x}}_{\rm{cur}},\bm{\sigma}_{\rm{cur}})$ $\triangleright$ Denoise sequence
10:  $\bm{d}\leftarrow(\bar{\bm{x}}_{\rm{cur}}-\hat{{\bm{y}}})/\bm{\sigma}_{\rm{cur}}$ $\triangleright$ Evaluate $\mathrm{d}\bar{\bm{x}}/\mathrm{d}{t}$ at $t_{\rm{cur}}$
11:  $\bar{\bm{x}}_{\rm{next}}\leftarrow\bar{\bm{x}}_{\rm{cur}}+(\bm{\sigma}_{\rm{next}}-\bm{\sigma}_{\rm{cur}})\bm{d}$ $\triangleright$ Euler step from $t_{\rm{cur}}$ to $t_{\rm{next}}$
12:  $n_{\rm{clean}}\leftarrow\lfloor t_{\rm{next}}\rfloor$ $\triangleright$ Infer finished snapshots. If 0, the following lines are a no-op
@@
15:  $\bar{\bm{x}}_{\rm{cur}}\leftarrow[\bar{\bm{x}}_{\rm{next}},{\bm{x}}_{\mathrm{new}}]$ $\triangleright$ Concatenate fresh noisy snapshots to futuremost
16:  $t_{\rm{cur}}\leftarrow t_{\rm{next}}-n_{\rm{clean}}$ $\triangleright$ Re-adjust global diffusion ‘time’ to be in $[0,1)$
@@ 5.1 Evaluation and Metrics
Due to the importance of probabilistic forecasts and uncertainty quantification, we focus on two key ensemble-based probabilistic metrics: Continuous Ranked Probability Score (CRPS) [matheson1976crps], and the spread-skill ratio (SSR) based on $M$-member ensembles, where $M=50$ for Navier-Stokes and $M=10$ for ERA5. The CRPS is a proper scoring rule commonly used to evaluate probabilistic forecasts [gneiting2014Probabilistic]. The spread-skill ratio is defined as the ratio of the square root of the ensemble variance to the corresponding ensemble-mean RMSE. It serves as a measure of the reliability of the ensemble, where values smaller than 1 indicate underdispersion, and larger values overdispersion [fortin2014ssr, rasp2023weatherbench2].
@@ 5.2 Navier-Stokes fluid dynamics
We use the Navier-Stokes fluid dynamics benchmark from [otness21nnbenchmark], defined on a $221\times 42$ grid. Each simulation features four randomly placed circular obstacles influencing the flow, with fluid viscosity set to $1\times 10^{-3}$. The dataset comprises $x$ and $y$ velocities and pressure fields. All models receive boundary conditions and obstacle masks as auxiliary inputs. For testing, models predict a 64-timestep trajectory from a single initial snapshot.
@@
We benchmark ERDM against DYffusion [cachay2023dyffusion], the current state-of-the-art method on this dataset, and PDE-Refiner [lippe2023pderefiner]. To ensure a fair comparison, we retrained DYffusion using our experimental setup, achieving an improvement of over $3\times$ on its originally reported CRPS scores. Our primary focus, however, is to evaluate ERDM’s performance relative to common EDM-derived approaches for dynamics forecasting. Consequently, our key baseline is an EDM denoiser parameterized by $D_{\theta}({\bm{x}}_{1}\ldots,{\bm{x}}_{W};\sigma,{\bm{y}}_{0})$, where all ${\bm{x}}_{w}$ are corrupted according to the same $\sigma$. With $W=1$ we recover a next-step forecasting conditional EDM as in [price2023gencast, kohl2023benchmarking]. We tried $W\in\{1,2,4,6\}$, and found $W=4$ to work best for EDM. All baselines are trained on three random seeds and share the same architecture as much as possible. For ERDM, we use $W=6,\sigma_{\text{min}}=0.002,\sigma_{\text{max}}=200,\rho=-10,P_{\text{mean}}=0.5,P_{\text{std}}=1.2,N=1.25$. Further details are provided in Section˜D.2.
@@ Results.
In Fig.˜4, we benchmark ERDM against the baselines on forecasting the five Navier-Stokes test trajectories of length 64. While the EDM-based baselines and PDE-Refiner start well in terms of CRPS, even beating ERDM for the very first few time steps, they exhibit a significantly higher error growth than ERDM. We attribute this to the explicit modeling of progressive uncertainty built into our model. At the end of the rollout, ERDM achieves a $50\%$ better CRPS than the best baseline, EDM with a window size of $W=4$. Similarly, DYffusion starts worse and catches up with the EDM $W=4$ baseline at the end of the rollout, but not with ERDM. In terms of calibration, we observe that ERDM consistently outperforms the EDM baselines, which are noticeably undercalibrated. We note that this is despite comprehensive tuning of these baselines, where we were only able to achieve calibration improvements at the significant expense of CRPS skill (not shown). Examples of generated forecasts are visualized as videos at this URL: https://youtu.be/jRcwZw5JLe0.
@@ 5.3 ERA5 weather forecasting
We benchmark our model on medium-range weather forecasting on the ERA5 reanalysis dataset. We use the $1.5^{\circ}$ resolution of the data (a $240\times 121$ grid) provided by [rasp2023weatherbench2], with 69 prognostic (input and output) variables: Temperature (t), geopotential (z), specific humidity (q), and the u and v components of wind (u,v) over 13 pressure levels (numbers after the abbreviation refer to the level in hPa) as well as the four surface variables 2m temperature (2t), mean sea level pressure (mslp), and u and v components of winds at 10m (10u, 10v). We train our models on $12$-hourly data, as in [price2023gencast], from 1979 to 2020 and evaluate on 64 2021 initial conditions at 00/12 UTC.
@@
Our primary baseline is a conditional next-step forecasting EDM, trained using a methodology consistent with ERDM’s development. To gauge the absolute performance of ERDM, we also benchmark it against two prominent external models: (1) IFS ENS: The European Centre for Medium-Range Weather Forecasts’ (ECMWF) operational, physics-based ensemble forecasting system [ifs-manual-cy46r1-ens]; (2) NeuralGCM ENS: A hybrid ML-physics stochastic model [kochkov2023neural]. Official $1.5^{\circ}$-resolution forecasts for both external models were sourced from Weatherbench-2 [rasp2023weatherbench2] for all variables where available. The IFS ENS evaluation uses the same 2021 initial condition dates as our EDM and ERDM models, but is verified against operational analyses. [^2] As opposed to ERA5 targets. This choice is standard practice [rasp2023weatherbench2] and favorable to IFS ENS. Due to the unavailability of NeuralGCM forecasts for 2021, we use its 2020 forecasts and evaluate them against the corresponding 2020 ERA5 data. Our conditional EDM baseline can be seen as a reproduction of GenCast [price2023gencast], for which forecasts are only available at higher spatial resolutions and a limited set of variables, but using our experimental setup and neural architecture. Notably, training ERDM is considerably less computationally expensive than NeuralGCM ENS $1.5^{\circ}$ (GenCast $1^{\circ}$), which required 128 (32) v5 TPUs and 10 (3.5) days to train. In contrast, ERDM was trained on 4 H200 GPUs in 5 days only. Lastly, we also include scores for z500 and 2t from Graph-EFM [oskarsson2024probabilistic], taken from their paper. These scores are provided for reference, noting that Graph-EFM is trained on 20 more years and evaluated in 2020. For ERDM, we use $W=6,\sigma_{\text{min}}=0.002,\sigma_{\text{max}}=500,\rho=-10,P_{\text{mean}}=2,P_{\text{std}}=1.2,N=2$. See Section˜D.2 for more experimental details.
Figure 6: Normalized spectral density of 14-day forecasts, averaged over high latitudes, $60^{\circ},90^{\circ}]$. Spectra are divided by the target ERA5 reanalysis spectra. ERDM generates highly accurate spectra that match or slightly beat the physics-based model IFS ENS. NeuralGCM ENS underestimates energy at the mid to high frequencies. See Section˜F.3 for more variables and absolute spectra.